Pittman Company is a small but growing manufacturer of telecommunications equipment. The company has no sales force of its owWeve already worked them up, said Barbara. Several companies we know about pay a 7.5% commission to their own salespeople4. Compute the degree of operating leverage that the company would expect to have at the end of next year assuming: a. The ag
Answer #1
Compute Pittman Cp's break Even Point in $ sales for
next Year assuming
a) the agen commission - Unchanged 15%
b) the agen commission - increased to 20%
c) compamy employed its own Sales force
Income Statement
Commission @15%$ Commission @20%$ Own Sales force$
Sales -a             2,35,00,000    2,35,00,000 2,35,00,000
Less
Manufacturing Cost             1,05,75,000    1,05,75,000 1,05,75,000
( Variable)
Sales Commission                35,25,000        47,00,000      17,62,500
( 15% on Revenue
( 20% on Revenue
( 7.5% on Revenue )
Total Variable cost-b             1,41,00,000    1,52,75,000 1,23,37,500
Contrbution margin(a-b)                94,00,000        82,25,000 1,11,62,500
Contrbution margin% 40% 35% 48%
Less
Fixed cost
Fixed Overhead                32,90,000        32,90,000      32,90,000
Fixed Marketing Expenses                   1,64,500          1,64,500      36,89,500 ( as below Cal)
Fixed Administratin Expenses                21,00,000        21,00,000      19,91,900
Fixed Interest cost                   8,22,500          8,22,500        8,22,500
Total Fixed cost                63,77,000        63,77,000      97,93,900
Income Before tax                30,23,000        18,48,000      13,68,600
Tax30%                   9,06,900          5,54,400        4,10,580
Net Profit after tax                21,16,100        12,93,600        9,58,020
Derived Fixed Marketing Expesnes - Own Sales force
Amnt($)
Fixed Marketing cost        1,64,500
Add- Amount pais as commission      35,25,000
Total Marketing cost      36,89,500
Derived Fixed Admin Expesnes - Own Sales force Amnt($)
Fixed admin Expenses ( as given in Question)      21,00,000
Less - Cost savings ( as mentioned in Question)        1,08,100
Fixed Admin Expenses      19,91,900
Calcuate $ Value Break Even Point
Commission @15%
Break Even Point Total Fixed cost/ contribution %
Break Even Point $ ($6377000/40%)    1,59,42,500
Commission @20%
Break Even Point Total Fixed cost/ contribution %
Break Even Point $ ($6377000/35%)    1,82,20,000
Commission Own Sales force
Break Even Point Total Fixed cost/ contribution %
Break Even Point $ ($9793900/48%)    2,06,18,737
Answer2 Need to calculate desired Sale to Achieve
As calculated above Amnt($)
Total Fixed cost                63,77,000
AddIncome Before tax                30,23,000
Total=A                94,00,000
Contribution margin-B 40%
Desired Sales $             2,35,00,000
Answer3 Determine volume of sales at which
net income would be Equal rgardless whether
Pitman company pay 20% commission or Own Sale force
At 20% commission Time
Contribution margin 35%
At Own Sales force
Contribution margin 48%
(100%-35%)*X+ Fixed cost = (100%-48%)*x+Fixed cost
65%*x+6377000= 52%*x+9793900
X represents- Volume of Required sales
65%x-52%x =9793900-6377000
x(65%-52%)= 3416900
x=             2,62,83,846
So Desired Sales Value required in Next Year $ 2,62,83,846